### v1

# NEXT: Multi-Grained Mixture of Experts via Text-Modulation for Multi-Modal Object Re-ID

Abstract
Multi-modal object re-identification (ReID) aims to extract identity features across heterogeneous spectral modalities to enable accurate recognition and retrieval in complex real-world scenarios. However, most existing methods rely on implicit feature fusion structures, making it difficult to model fine-grained recognition strategies under varying challenging conditions. Benefiting from the powerful semantic understanding capabilities of Multi-modal Large Language Models (MLLMs), the visual appearance of an object can be effectively translated into descriptive text. In this paper, we propose a reliable multi-modal caption generation method based on attribute confidence, which significantly reduces the unknown recognition rate of MLLMs in multi-modal semantic generation and improves the quality of generated text. Additionally, we propose a novel ReID framework NEXT, the Multi-grained Mixture of Experts via Text-Modulation for Multi-modal Object Re-Identification. Specifically, we decouple the recognition problem into semantic and structural expert branches to separately capture modality-specific appearance and intrinsic structure. For semantic recognition, we propose the Text-Modulated Semantic-sampling Experts (TMSE), which leverages randomly sampled high-quality semantic texts to modulate expert-specific sampling of multi-modal features and mining intra-modality fine-grained semantic cues. Then, to recognize coarse-grained structure features, we propose the Context-Shared Structure-aware Experts (CSSE) that focuses on capturing the holistic object structure across modalities and maintains inter-modality structural consistency through a soft routing mechanism. Finally, we propose the Multi-Modal Feature Aggregation (MMFA), which adopts a unified feature fusion strategy to simply and effectively integrate semantic and structural expert outputs into the final identity representations. Extensive experiments on four common public datasets demonstrate the effectiveness of our method and show that it significantly outperforms existing state-of-the-art methods. The source code will be released upon acceptance.
1 Introduction
Figure 1: Motivation of our proposed method. Better view with color and zoom in.
Figure 2: Generated caption samples and text quality analysis on RGBNT201 dataset.
Object re-identification (ReID) aims to construct discriminative identity representations by capturing object characteristics such as appearance, clothing, body shape, and posture [1, 2, 3, 4, 5, 6, 7]. However, in real-world scenarios, variations in lighting, weather, occlusion, and clothing [8, 9, 10, 11] hinder the performance of conventional ReID methods. To address these challenges, multi-modal object re-identification [12, 13, 14, 15] has attracted increasing research interest in recent years. Benefit from the complementary advantages of different spectral modalities, identity information can be more comprehensively captured. Despite this, intra-modal noise and inter-modal heterogeneity remain key obstacles to effective identity representation fusion. To overcome these issues, existing methods focus on implicit identity-level feature fusion [16, 17, 18, 19], fine-grained feature complementarity [20, 21, 15], and frequency-domain interactions [22, 20]. However, semantic-aware fusion [23, 24] of multi-modal identity features is still underexplored.
Recent progress in vision-language models [25, 26, 27, 28, 29] has advanced visual understanding tasks, including object re-identification [30, 31]. Traditional ReID methods benefit from the latent alignment space between visual and linguistic modalities, enabling effective extraction of semantic identity representations [30, 31]. Recent research on learnable semantic prompts allows models to capture identity-relevant signals even in the absence of explicit semantic descriptions [32, 33, 24]. With the evolution of Multi-modal Large Language Models (MLLMs) [34, 35, 36], the ability to generate image-text pairs introduces new opportunities for object re-identification. Pioneering methods such as IDEA [23], TVI-LFM [37], and MP-ReID [38] integrate MLLMs into the ReID task by generating identity-relevant captions, thereby extending identity representations into the textual modality. For instance, IDEA [23] invests textual features into the visual representation space to capture cross-modal semantic information. However, the perception and fusion of fine-grained identity semantics across multi-modal data remain underexplored. Therefore, we raise the question: Can we further modulate the model to perceive fine-grained multi-modal features under diverse identity semantics?
Compared to labor-intensive human annotation, MLLMs can easily obtain textual descriptions of objects using instruction-based templates. However, in multi-modal scenarios, the appearance information of the object is often suppressed or distorted by various environmental conditions and noise. The IDEA [23] introduces a caption generation method combined with attribute-based refinement to leverage modal-specific descriptions. Nevertheless, it overlooks the accuracy of captions in low-quality modalities, leading to unreliable quantification of attributes. To tackle this, we propose a reliable caption generation framework, as shown in Fig. 1(c). Specifically, we decompose the caption generation into two steps: confidence-aware attribute generation and multi-modal attributes aggregation. By enforcing the model to output a confidence score for each attribute, we encourage fine-grained evaluation and strengthen the expressiveness of each attribute. Based on the confidence scores, we quantify the attribute importance across different modalities, enabling us to complement missing semantic information through multi-modal captions. Finally, the aggregated attributes are passed to an LLM to flexibly generate the final textual captions.
High-quality captions effectively guide the model to focus on identity-relevant representations. Building on this, we propose a multi-grained mixture of experts framework named NEXT, which models the identity recognition process through semantic-sampling and structure-aware experts. The framework consists of three main components: Text-Modulated Semantic-sampling Experts (TMSE), Context-Shared Structure-aware Experts (CSSE), and Multi-Modal Feature Aggregation (MMFA). Specifically, TMSE samples fine-grained part-level features within each modality. The text modulation mechanism encourages the experts to focus on semantically relevant regions during training, thereby capturing intra-modality identity semantic features. Meanwhile, CSSE perceives inter-modality structural features of the object identity through a soft routing mechanism shared across modalities. Finally, MMFA unifies the features of different expert branches into a comprehensive multi-modal identity representation. Extensive experiments on four multi-modal object re-identification benchmarks demonstrate the effectiveness of our proposed approach.
The main contributions of this paper are summarized as follows:
  • We propose a novel multi-expert learning framework to learn comprehensive multi-modal identity representations for multi-modal object re-identification. It decomposes the identity modeling process into semantic and structural recognition to better leverage the complementary information across heterogeneous modalities.
  • We propose the Text-Modulated Semantic-sampling Experts (TMSE) to model intra-modal feature sampling under diverse semantic challenges, the Context-Shared Structure-aware Experts (CSSE) to capture the complete inter-modal identity structure, and the Multi-Modal Feature Aggregation (MMFA) to integrate the outputs of multiple experts into a unified identity representation.
  • We propose a novel and reliable multi-modal caption generation method that leverages attribute confidence to effectively harness MLLMs to expand four popular benchmarks with high-quality textual annotations. Extensive experiments are conducted on these benchmarks to validate the effectiveness of our method. The results demonstrated that the proposed method significantly outperformed the state-of-the-art methods.
2 Related Work
2.1 Multi-Modal Object Re-Identification
Multi-modal object ReID leverages the complementary imaging advantages of multi-spectra to enable robust identity recognition under adverse conditions. However, intra-modal noise and inter-modal spectral heterogeneity remain significant challenges. To fuse the heterogeneous visual features, TOP-ReID [16] proposes a modality permutation and reconstruction mechanism to align multi-modal representations. PromptMA [17] introduces a prompt-based token selection and fusion strategy to effectively integrate multi-modal features and suppress background interference. MambaPro [19] presents a selective state-space fusion approach based on Mamba [39] to aggregate intra- and inter-modal features. EDITOR [20] adopts frequency- and feature-based selection mechanisms to filter out background and low-quality noise. FACENet [15] utilizes illumination priors to enhance degraded modalities. TIENet [22] introduces an amplitude-guided phase learning strategy for frequency-domain enhancement. ICPL-ReID [24] applies an identity prototype-based conditional prompt learning method to guide semantic learning across modalities. IDEA [23] inverts textual features into the visual space to guide model learning, and captures discriminative local features across multi-modal via deformable aggregation. Despite these advancements, existing methods still lack a deep exploration of identity semantics and intrinsic structures. To bridge this gap, we propose a multi-grained mixture of experts via text-modulation that fuses features from different modalities through semantic and structural perspectives, aiming to enhance multi-modal identity representation.
2.2 Large Foundation Model in Re-Identification
The rise of MLLMs [35, 34, 36] significantly advances visual understanding tasks. Numerous studies demonstrate the effectiveness of MLLMs [40, 41] in generating high-quality text and robust generalization in vision-language alignment. Large-scale vision-language foundation models, such as CLIP [25], exhibit powerful visual perception and semantic understanding capabilities. Leveraging the latent image-text aligned space, researchers employ the learnable prompt [29, 28] to effectively adapt these models to classification, retrieval, and open-set recognition tasks. Methods like CLIP-ReID [30] and PromptSG [31] utilize learnable semantic prompts to transfer the CLIP [25] model to object re-identification tasks. Other approaches, such as MP-ReID [38], TVI-LFM [37], and IDEA [23], exploit MLLMs to generate identity-level textual descriptions from existing ReID datasets. However, these methods still lack reliable text generation under multi-modal conditions. Unlike IDEA [23], we further quantify attributes and propose a confidence-aware attribute generation strategy. We incorporate a multi-modal merging and complementation mechanism to achieve high-quality and reliable caption generation. Additionally, we introduce the NEXT framework, which leverages semantic experts to extract fine-grained multi-modal part features, aiming to tackle diverse semantic challenges in multi-modal scenarios.
2.3 Multi-Modal Mixture of Experts
Mixture-of-Experts (MoE) [42] achieves remarkable progress in both computer vision and natural language processing domains due to its flexible design and efficient activation mechanisms. In multi-modal visual tasks [43, 44, 45, 46], MoE frameworks are typically designed with modality-shared and modality-specific experts to effectively capture diverse multi-modal features. However, these methods still fall short in learning identity-related representations for multi-modal object ReID. DeMo [18] first applies the MoE framework to the multi-modal ReID task by combining modality-specific experts to decouple shared and unique features. Despite this, it lacks fine-grained identity semantic understanding. To address this limitation, we propose text-modulated semantic-sampling experts and context-shared structure-aware experts to jointly capture semantic and structural features, enabling more comprehensive identity representation across multi-modalities.
3 Multi-Grained Mixture of Experts via Text-Modulation
In this section, we elaborate the proposed mixture-of-experts learning framework. As shown in Fig. 3 our proposed Multi-Grained Mixture-of-Experts framework includes the Text-Modulated Semantic-sampling Experts (TMSE) module, Context-Shared Structure-aware Experts (CSSE) module and Multi-Modal Feature Aggregation (MMFA) module.
Overview. We first define the basic notations and introduce the overall pipeline of the proposed framework. For multi-modal object ReID, each identity instance consists of three different modalities: visible light (RGB), near-infrared (NIR), and thermal infrared (TIR), denoted as $\textbf{X}_{I}=[X_{I,\mathrm{rgb}},X_{I,\mathrm{nir}},X_{I,\mathrm{tir}}]$. Benefiting from the strong visual understanding capabilities of Multi-modal Large Language Models (MLLMs), we input each modality image $\textbf{X}_{I}$ along with the modality-specific instruction template $\textbf{P}_{M}$ into the MLLM $\mathcal{M}(\cdot)$ to extract detailed identity appearance information $\textbf{X}_{T}=[X_{T,\mathrm{rgb}},X_{T,\mathrm{nir}},X_{T,\mathrm{tir}}]$. Supported by large-scale pretraining for vision-text alignment, CLIP effectively aligns image and text features within a shared latent space. We feed the visual modality $\textbf{X}_{I}$ into the CLIP visual encoder $\mathcal{V}(\cdot)$ to extract visual features $\mathbf{F}_{I,m}=[\mathbf{f}_{I,m}^{\mathrm{cls}};\mathbf{F}_{I,m}^{\mathrm{tok}}]\in\mathbb{R}^{M\times(1+N)\times D}$, and input the object appearance description $\textbf{X}_{T}$ into the frozen CLIP text encoder $\mathcal{T}(\cdot)$ to extract textual features $\mathbf{F}_{T,m}=[\mathbf{f}_{T,m}^{\mathrm{cls}};\mathbf{F}_{T,m}^{\mathrm{tok}}]\in\mathbb{R}^{M\times(1+L)\times D}$, where $M$ is modality number, $N$ is the length of visual patch tokens, $L$ is the length of text patch tokens, $D$ is the dimension length. As shown in Fig. 3(b), we define a set of semantic-sampling experts $\textbf{E}_{T}=\{E_{t}^{(1)},E_{t}^{(2)},\dots,E_{t}^{(N_{T})}\}$, where $N_{T}$ denotes the number of semantic experts. Guided by high-quality textual features, each expert discards semantically irrelevant tokens and mines semantically related patch tokens within each modality. Meanwhile, as depicted in Fig. 3(d), we introduce another set of structure-aware experts $\textbf{E}_{C}=\{E_{c}^{(1)},E_{c}^{(2)},\dots,E_{c}^{(N_{C})}\}$, where $N_{C}$ denotes the number of structure experts. These experts jointly utilize all patch features and adopt a soft routing mechanism to perceive the structural consistency of objects across modalities. Finally, we aggregate the features captured by all experts, including $\textbf{E}_{T}$ and $\textbf{E}_{C}$, through the multi-modal feature aggregation module to form a unified multi-modal representation as the final discriminative feature.
Figure 3: Framework of our proposed method.
3.1 Text-Modulated Semantic-sampling Experts (TMSE)
In practical scenarios, challenges such as illumination variations, background interference, and low-quality noise often affect different local regions across modalities. This necessitates overcoming modality heterogeneity and discovering complementary cross-modal features. However, existing approaches often lack explicit semantic guidance. To address this, we leverage the object-level textual semantics $\mathbf{X}_{T}$ extracted from MLLMs to guide a set of semantic experts $\textbf{E}_{T}=\{E_{t}^{(1)},E_{t}^{(2)},\dots,E_{t}^{(N_{T})}\}$ in dynamically sampling semantic-related local features across different modalities, to tackle various semantic challenges. The structure of each expert is defined as:
$E_{t}(\mathbf{F})=\mathrm{Dropout}\left(\mathrm{MLP}\left(\mathrm{LN}(\mathbf{F})\right)\right)+\mathbf{F},$ (1)
where $\mathbf{F}$ denotes the input features, $\mathrm{LN}$ is a layer normalization, $\mathrm{MLP}$ is a multilayer perceptron, and $\mathrm{Dropout}$ is applied to prevent overfitting.
Dynamic Sampling Route. To sample semantic features from each modality, we design a set of modality-specific sampling routes for each expert, denoted as $\mathbf{R}_{m}=[\mathbf{R}_{rgb},\mathbf{R}_{nir},\mathbf{R}_{tir}]$. As shown in Eq.(2), we encode the patch-level features $\mathbf{F}_{I,m}^{\mathrm{tok}}$ into the routing matrix $\bm{\alpha}_{m}$, and map the class token $\mathbf{f}_{I,m}^{\mathrm{cls}}$ into the scalar threshold $\bm{\sigma}_{m}$. The formulation is as follows:
$\bm{\alpha}_{m}=\mathrm{FC}\left(\delta\left(\mathrm{FC}(\mathbf{F}_{I,m}^{\mathrm{tok}})\right)\right),\quad\bm{\sigma}_{m}=\mathrm{FC}\left(\delta\left(\mathrm{FC}(\mathbf{f}_{I,m}^{\mathrm{cls}})\right)\right),\quad\bm{\alpha}_{m}\in\mathbb{R}^{H\times W},\bm{\sigma}_{m}\in\mathbb{R}.$ (2)
$\mathbf{M}_{m}(i,j)=\begin{cases}1,&\text{if }\bm{\alpha}_{m}(i,j)>\bm{\sigma}_{m}\\ 0,&\text{otherwise}\end{cases},\quad\text{for }i=1,\dots,H,\ j=1,\dots,W$ (3)
[…]
Multi-Grained Text Modulation. To empower the semantic awareness of each expert, we employ random semantic text prompts during training, guiding each expert toward distinct semantic challenges. Formally, the raw semantic description is decomposed into fine-grained sentences for each modality: $X_{T,m}=\{s_{m}^{(1)},s_{m}^{(2)},\dots,s_{m}^{(n_{m})}\},\quad m\in\{\mathrm{rgb},\mathrm{nir},\mathrm{tir}\}$, where $n_{m}$ is the maximum number of sentences. For each expert $E_{t}^{(i)}$, we randomly select a subset $\hat{X}_{T,m}^{(i)}$ from $X_{T,m}$ as its semantic modulation signal. The sampling process is defined as:
$\hat{\mathbf{X}}_{T}^{(i)}=[\hat{X}_{T,rgb}^{(i)},\hat{X}_{T,nir}^{(i)},\hat{X}_{T,tir}^{(i)}],\quad\hat{X}_{T,m}^{(i)}\sim\mathbf{Sampling}(X_{T,m})$ (4)
where $\mathrm{i}$ denotes the $\mathrm{i}$-th semantic experts, and $\mathbf{Sampling}$ is a random probability distribution.
To modulate the visual sampling route, we first encode the randomly sampled high-quality semantic text into the visual-text latent space as text feature vector $\mathbf{f}^{\mathrm{cls}}_{\hat{T},m}$. Then, we compute the cosine similarity between this text vector and the visual patch tokens to obtain the semantic relevance matrix $\bm{\beta}_{m}$:
$\bm{\beta}_{m}=\frac{\mathbf{f}^{\mathrm{cls}}_{\hat{T},m}\cdot\mathbf{F}^{\mathrm{tok}}_{I,m}}{\|\mathbf{f}^{\mathrm{cls}}_{\hat{T},m}\|_{2}\cdot\|\mathbf{F}^{\mathrm{tok}}_{I,m}\|_{2}},\quad\bm{\beta}_{m}\in\mathbb{R}^{H\times W}.$ (5)
We then integrate this semantic guidance with the route matrix $\bm{\alpha}_{m}$ to produce the modulation matrix $\bm{\gamma}_{m}$, as formulated below:
$\bm{\gamma}_{m}=\mathbf{Modulation}\left(\bm{\alpha}_{m},\bm{\beta}_{m}\right)+\bm{\alpha}_{m},\quad\bm{\gamma}_{m}\in\mathbb{R}^{H\times W}$ (6)
where $\mathbf{Modulation}$ refers to the modulation network whose structure is illustrated in Fig. 3(c).
Finally, we replace the route matrix $\bm{\alpha}_{m}$ in Eq.(3) with the modulated matrix $\bm{\gamma}_{m}$ to obtain the final modulation-based sampling matrix $\hat{\mathbf{M}}_{m}$.
$\hat{\mathbf{F}}_{I}=\mathrm{Concat}\left(\hat{\mathbf{M}}_{m}\odot\hat{\mathbf{F}}_{I,m}^{\mathrm{tok}}\right),\quad\hat{\mathbf{F}}_{I,m}^{\mathrm{tok}}=E_{t}(\mathbf{F}_{I,m}^{\mathrm{tok}}),\quad\forall m\in\{\mathrm{rgb},\mathrm{nir},\mathrm{tir}\}$ (7)
As shown in the Eq.(7) above, $\hat{\mathbf{M}}_{m}$ dynamically samples the informative features for each semantic expert $\mathrm{E}^{(i)}_{t}$ within a modality. These sampled features are concatenated to obtain the final multi-modal semantic feature $\hat{\mathbf{F}}_{I}$.
3.2 Context-Shared Structure-aware Experts (CSSE)
Although semantic experts focus on fine-grained part sampling under various semantic challenges. The identity recognition and discrimination still rely on the structural integrity of the object. To this end, we introduce a set of structure experts, denoted as $\mathbf{E}_{C}=\{E_{c}^{(1)},E_{c}^{(2)},\dots,E_{c}^{(N_{C})}\}$, to comprehensively perceive the holistic object structure across modalities.
Specifically, to preserve structural integrity, we first concatenate multi-modal patch features as $\mathbf{F}_{I}^{\mathrm{tok}}=\mathrm{Concat}\left(\mathbf{F}_{I,\mathrm{rgb}}^{\mathrm{tok}},\mathbf{F}_{I,\mathrm{nir}}^{\mathrm{tok}},\mathbf{F}_{I,\mathrm{tir}}^{\mathrm{tok}}\right)$, and then feed them into a modality-shared routing network $R_{s}$ to compute the expert fusion weights as follows:
$\bm{\omega}=\mathbf{Softmax}\left(\mathrm{FC}\left(\mathbf{F}_{I}^{\text{tok}}\right)\right),\quad\bm{\omega}\in\mathbb{R}^{N_{C}}$ (8)
where $\mathrm{FC}$ denotes the fully connected layer, and $\mathbf{Softmax}$ encodes the expert selection weights, resulting in the soft routing matrix $\bm{\omega}$.
Finally, we feed the concatenated multi-modal token features into the structure experts module $\mathrm{E}_{C}$ and compute the final multi-modal structural representation by weighting each output of expert using the soft routing matrix $\bm{\omega}$. The process is formulated as follows:
$\tilde{\mathbf{F}}_{I}=\sum_{i=1}^{N_{C}}\omega_{i}\cdot E^{(i)}_{c}(\mathbf{F}_{I}^{\text{tok}}),\quad\tilde{\mathbf{F}}_{I}^{tok}=$$\{E^{(i)}_{c}(\mathbf{F}_{I}^{\text{tok}})\mid i=1,2,\dots,N_{C}\},$ (9)
where $\tilde{\mathbf{F}}_{I}$ is represent the final multi-modal structure feature.
3.3 Multi-Modal Feature Aggregation (MMFA)
Experts of different types mine various identity features. The semantic expert $\mathrm{E}_{T}$ extracts fine-grained semantic features $\hat{\mathbf{F}}_{I}^{(i)}$, and the structural expert $\mathrm{E}_{C}$ captures structural consistency features $\tilde{\mathbf{F}}_{I}$. As illustrated in Fig. 3(e), we combine these features to form the complete object identity feature set $\mathbf{F}_{I}^{\mathrm{exp}}=[\hat{\mathbf{F}}_{I}^{(1)},\hat{\mathbf{F}}_{I}^{(2)},\dots,\hat{\mathbf{F}}_{I}^{(N_{T})};\tilde{\mathbf{F}}_{I}]$. The modality class tokens are concatenated to form the modality query features $\mathbf{f}_{I}^{\mathrm{cls}}=\mathrm{Concat}\left(\mathbf{f}_{I,\mathrm{rgb}}^{\mathrm{cls}},\mathbf{f}_{I,\mathrm{nir}}^{\mathrm{cls}},\mathbf{f}_{I,\mathrm{tir}}^{\mathrm{cls}}\right)$. Following this, we treat $\mathbf{f}_{I}^{\mathrm{cls}}$ as the query feature Q and the expert features as the key K and value V. Through the Cross-Attention mechanism, we obtain the multi-modal representation for each expert and concatenate them to form the final identity representation $\hat{\mathbf{f}}_{I}^{\mathrm{cls}}$. The process is formulated as follows:
$\hat{\mathbf{f}}_{I}^{\mathrm{cls}}=\mathrm{Concat}\left(\hat{\mathbf{f}}_{I}^{\mathrm{cls(1)}},\hat{\mathbf{f}}_{I}^{\mathrm{cls(2)}},\dots,\hat{\mathbf{f}}_{I}^{\mathrm{cls(N_{T})}};\tilde{\mathbf{f}}_{I}^{\mathrm{cls}}\right),$ (10)
$\hat{\mathbf{f}}_{I}^{\mathrm{cls(i)}}=\mathbf{FFN}\left(\mathrm{LN}\left(\mathbf{CA}\left(\mathbf{f}_{I}^{\mathrm{cls}},\mathbf{F}_{I}^{\mathrm{exp(i)}}\right)\right)\right),\quad\text{for }i=1,\dots,N_{T}+1.$ (11)
where $\mathbf{CA}$ is the Cross-Attention mechanism, and $\mathbf{FFN}$ is the feedback forward layer.
3.4 Optimization and Inference
In line with prior work [3, 5], we use the object identity ID label as the ground-truth to train the classification loss $\textit{L}_{\mathrm{id}}$, and adopt the triplet loss $\textit{L}_{\mathrm{tri}}$ to enhance identity compactness and separability. The final loss function is defined as:
$\textit{L}_{\mathrm{final}}(\hat{\mathbf{f}}_{I}^{\mathrm{cls}})=\textit{L}_{\mathrm{id}}(\hat{\mathbf{f}}_{I}^{\mathrm{cls}})+\textit{L}_{\mathrm{tri}}(\hat{\mathbf{f}}_{I}^{\mathrm{cls}}).$ (12)
Table 1: Comparison with the state-of-the-art methods on RGBNT201. The best and second best results are marked in bold and underline, respectively.
Methods Venue Structure RGBNT201
mAP R-1 R-5 R-10
HAMNet [12] AAAI20 CNN 27.7 26.3 41.5 51.7
PFNet [14] AAAI21 CNN 38.5 38.9 52.0 58.4
IEEE [47] AAAI22 CNN 46.4 47.1 58.5 64.2
TIENet [22] TNNLS25 CNN 54.4 54.4 66.3 71.1
UniCat [48] NIPSW23 CNN 57.0 55.7 - -
HTT [49] AAAI24 ViT 71.1 73.4 83.1 87.3
TOP-ReID [16] AAAI24 ViT 72.3 76.6 84.7 89.4
EDITOR [20] CVPR24 ViT 66.5 68.3 81.1 88.2
WTSF-ReID [50] ESWA25 ViT 67.9 72.2 83.4 89.7
ICPL-ReID [24] TMM25 CLIP 75.1 77.4 84.2 87.9
PromptMA [17] TIP25 CLIP 78.4 80.9 87.0 88.9
MambaPro [19] AAAI25 CLIP 78.9 83.4 89.8 91.9
DeMo [18] AAAI25 CLIP 79.0 82.3 88.8 92.0
IDEA [23] CVPR25 CLIP 80.2 82.1 90.0 93.3
NEXT Ours CLIP 82.4 86.6 92.0 94.7
4 Experiment
4.1 Datasets and Evaluation Protocols
Datasets. We evaluate our method on four multi-modal object ReID datasets: RGBNT201 [14], MSVR310 [13], RGBNT100 [12], and WMVEID863 [15]. To extend these datasets, we employ GPT-4o [34] and Qwen-VL [35] to automatically generate object attribute with confidence, and DeepSeek-V3 [51] to compose the final caption for each image modality in the train and test sets. Details of dataset scales and implementation details are provided in the Appendix A.
Evaluation Protocols. In line with the convention of the ReID community [3, 1], we use Rank-K (K = 1, 5, 10) matching accuracy and the Mean Average Precision (mAP) as evaluation metrics. As in previous works [14, 12, 15], we adopt the common evaluation protocol for RGBNT201 [14], RGBNT100 [12] and WMVEID863 [15]. For MSVR310 [13], we enforce a strict protocol [13] that filters out samples with the same identity and time span based on time labels to avoid easy matching.
4.2 Comparison with State-of-the-Art Methods
Performance on Person Dataset. In Table 1, we compare our NEXT with existing multi-modal methods on RGBNT201 [14] dataset. Benefiting from the flexible fusion of semantic and structure experts, our method achieves a significant performance lead, reaching 82.4%/86.6% mAP/Rank-1 accuracy. Compared with DeMo [18], which adopts MoE structure to handle modality-shared and modality-specific features, our method demonstrates superior performance by perceiving both the semantic and structural features of the objects. Against IDEA [23], which utilizes semantic inversion and deformable offset sampling, our method leverages text-modulation to guide expert sampling multi-modal semantic, improving mAP and Rank-1 by +2.2% and +4.5%, respectively. These results validate the effectiveness of NEXT in learning more discriminative multi-modal identity features through textual semantics.
Table 2: Comparison with the state-of-the-art methods on MSVR310, RGBNT100 and WMVEID863.
Methods Venue Strucutre MSVR310 RGBNT100 WMVEID863
mAP R-1 mAP R-1 mAP R-1 R-5 R-10
HAMNet [12] AAAI20 CNN 27.1 42.3 74.5 93.3 45.6 48.5 63.1 68.8
PFNet [14] AAAI21 CNN 23.5 37.4 68.1 94.1 50.1 55.9 68.7 75.1
IEEE [47] AAAI22 CNN 21.0 41.0 61.3 87.8 45.9 48.6 64.3 67.9
CCNet [13] INFFUS23 CNN 36.4 55.2 77.2 96.3 50.3 52.7 69.6 75.1
TOP-ReID [16] AAAI24 ViT 35.9 44.6 81.2 96.4 67.7 75.3 80.8 83.5
FACENet [15] INFFUS25 ViT 36.2 54.1 81.5 96.9 69.8 77.0 81.0 84.2
EDITOR [20] CVPR24 ViT 39.0 49.3 82.1 96.4 65.6 73.8 80.0 82.3
MambaPro [19] AAAI25 CLIP 47.0 56.5 83.9 94.7 69.5 76.9 80.6 83.8
PromptMA [17] TIP25 CLIP 55.2 64.5 85.3 97.4 - - - -
DeMo [18] AAAI25 CLIP 49.2 59.8 86.2 97.6 68.8 77.2 81.5 83.8
IDEA [23] CVPR25 CLIP 47.0 62.4 87.2 96.5 - - - -
ICPL-ReID [24] TMM25 CLIP 56.9 77.7 87.0 98.6 67.2 74.0 81.3 85.6
NEXT Ours CLIP 60.8 79.0 88.2 97.7 70.9 77.8 84.3 86.7
Table 3: Ablation studies of different modules.
Modules Metrics
MMFA TMSE CSSE mAP R-1
A 71.0 73.6
B 74.2 76.6
C 78.9 84.4
D 82.4 86.6
Table 4: Effectiveness of different route strategy.
Router Type Metrics
$\mathbf{E}_{T}$ $\mathbf{E}_{C}$ mAP R-1
A 78.3 82.5
B 78.7 83.4
C 80.3 84.3
D 82.4 86.6
Table 5: Effectiveness of caption quality.
Methods Quality mAP R-1
IDEA 35% 76.1 77.9
70% 78.1 79.9
100% 80.2 84.0
Ours 35% 77.1 79.7
70% 80.0 82.2
100% 82.4 86.6
Figure 4: Effectiveness of the number of experts.
Performance on Vehicle Dataset. Table 2 outlines the performance of our method on vehicle ReID datasets. On the MSVR310 [13] dataset, our approach significantly outperforms existing methods, achieving 60.8%/79.0% mAP/Rank-1 accuracy, attributed to its semantic sampling and structural perception. On the RGBNT100 [12] dataset, our method continues to lead in mAP performance. For the challenging WMVEID863 dataset, where lighting variations degrade recognition performance, our method still improves over FACENet by +1.1%/+0.8% mAP/Rank-1. These results confirm the strong generalization and robustness of NEXT in vehicle ReID under complex conditions.
4.3 Ablation Studies
On the RGBNT201 [14] dataset, we thoroughly evaluation the effectiveness of each proposed module. Starting from a baseline built on a three-branch CLIP [25] visual backbone, we incrementally incorporate our components and remove them individually to investigate their specific contributions to the model’s overall performance. More experiments are available in the Appendix B.
Effectiveness of Key Modules. As shown in Table 7, the baseline (Row A) achieves 71.0% mAP and 73.6% Rank-1. In Row B, the MMFA improves performance by +3.2%/+3.0% by fusing baseline features as a single expert. Adding the TMSE module in Row C boosts accuracy to 78.9% mAP and 84.4% Rank-1 under semantic modulation. With the CSSE in Row D, the model gains performance +2.5%/+2.2% in mAP/Rank-1. These results confirm the contribution of each module to the overall performance.
Effectiveness of Route Strategy. As shown in Table 4, using modality-specific routing for both experts (Row A) results in 78.3%/82.5% mAP/Rank-1. Sharing routing for the semantic experts (Row B) slightly improves performance to 78.7%/83.4%. When the structural experts adopt modality-shared routing (Row C), the model gains an additional +1.6%/+0.9% In Row D, reverting the semantic expert to modality-specific routing further improves the results, showing that intra-modal semantic focus helps tackle semantic challenges more effectively.
Effectiveness of Caption Quality. In Table 5, we evaluate how text quality affects model performance by corrupting the captions with varying levels of noise. At 100% text quality, all models reach optimal accuracy, with IDEA achieving its best Rank-1 score. However, as quality degrades to 70%, performance drops significantly. When only 35% of the original text remains, IDEA falls to 76.1%/77.9% (mAP/Rank-1), and NEXT to 77.1%/79.7%. This highlights the strong dependency of semantic-guided methods on high-quality text.
Effectiveness of Experts Number. Fig. 4 illustrates how varying the number of experts influences model accuracy. With few experts, the model struggles to learn rich semantic information or recognize structural identity cues, which hampers overall performance. Conversely, an excessive number of experts causes redundancy and ambiguity in semantic-structural fusion, impairing identity representation. The results suggest that an appropriately chosen number of experts offers the best trade-off and leads to optimal performance.
Figure 5: Visualization of the sampled patch tokens and activated feature regions for experts.
4.4 Visualization Analysis
Expert Feature Visualization. As shown in Fig. 5, we visualize the sampling masks of the semantic experts and the activation maps of the structural experts. We observe that different semantic experts $E^{(i)}_{t}$ exhibit distinct modality preferences: for instance, $E^{(1)}_{t}$ tends to sample features from the RGB modality, $E^{(2)}_{t}$ from the NIR modality, and $E^{(3)}_{t}$ from the TIR modality. These sampling patterns are notably complementary across modalities, indicating that the semantically modulated expert network effectively learns complementary modality-specific features to enhance unimodal representations. Structural experts, on the other hand, maintain the integrity of coarse-grained identity structures. The parts ignored by expert $E^{(3)}_{c}$ are well captured by other structural experts. These observations validate that NEXT successfully decouples identity recognition into semantic sampling and structure modeling. More visualization results are provided in the Appendix D.
5 Conclusion
In this paper, we focus on the semantic learning of multi-modal object ReID. We first propose a reliable multi-modal caption generation approach based on MLLMs, which use confidence-aware attribute prompts to construct structured attribute sets, and quantitatively merge multi-modal attributes to compose high-quality captions. Additionally, we propose the NEXT, a novel multi-grained MoE network for multi-modal ReID. We decouple the recognition problem into fine-grained semantic sampling and coarse-grained structural perception to overcome modality heterogeneity and achieve discriminative identity representations. Specifically, we propose Text-Modulated Semantic-sampling Experts (TMSE) to perceive fine-grained identity semantics and dynamically sample semantic parts. Then, we propose Context-Shared Structure-aware Experts (CSSE) to capture coarse-grained structural cues via modality-shared routing. Finally, we incorporate a Multi-Modal Feature Aggregation (MMFA) module to unify features from various experts and obtain complete identity representations. Extensive experiments on four popular benchmarks demonstrate the effectiveness of our method.
Appendix A Datasets Scale and Implementation Details
A.1 Datasets Scale
Table 6: Statistical analysis of datasets used in our experiments.
Datasets Object Type # Samples # IDs # Cams
RGBNT201 [14] Person 4,787 201 4
MSVR310 [13] Vehicle 2,087 310 8
RGBNT100 [12] Vehicle 17,250 100 8
WMVEID863 [15] Vehicle 4,709 863 8
We list the dataset statistics in Table 6, including the number of samples, identities, and cameras across these datasets. RGBNT201 [14] focuses on person ReID with four non-overlapping camera views and samples composed of RGB, NIR, and TIR modalities. MSVR310 [13] captures vehicles from eight unique angles over extended durations. RGBNT100 [12] provides diverse vehicle pairs across three spectra. WMVEID863 [15] serves as the largest available tri-spectral vehicle dataset with harsh lighting challenges.
A.2 Implementation Details
We resize each spectral image to $256\times 128$ for person samples, and $128\times 256$ for vehicle samples. Data augmentation includes random horizontal flipping, padding, cropping, and erasing [52]. We utilize the vision and text encoders in CLIP as the backbone, freezing all parameters of the text branch for text semantic projection, while keeping the visual branch fully trainable. Adam optimizer is used with the learning rate of 3.5e-6, weight decay of 0.0001, and momentum set to 0.9. All experiments are conducted on one NVIDIA RTX 4090 GPU using the PyTorch framework.
Appendix B Ablation Studies
Table 7: Ablation studies of different modules on MSVR310 and RGBNT201.
Modules MSVR310 RGBNT201
MMFA TMSE CSSE mAP R-1 R-5 R-10 mAP R-1 R-5 R-10
A 50.1 68.9 83.4 87.8 71.0 73.6 84.2 88.2
B 55.0 73.3 86.8 91.7 74.2 76.6 87.8 91.3
C 59.7 75.1 88.2 92.0 78.9 84.4 91.1 93.2
D 60.8 79.0 89.2 92.2 82.4 86.6 92.0 94.7
B.1 Effectiveness of Key Modules.
As shown in Table 7, to thoroughly validate the effectiveness of each proposed module, we conduct ablation studies on the vehicle dataset MSVR310 [13]. We also provide the complete ablation results on the RGBNT201 [14] dataset. In Row A, we use a three-branch CLIP visual encoder as the baseline, achieving 50.1% mAP and 68.9% Rank-1 accuracy on the vehicle dataset. By introducing the MMFA module, which treats each backbone feature as an independent expert, the performance improves to 55.0% mAP and 73.3% Rank-1. Further adding TMSE semantic experts brings an additional +4.7% mAP and +1.8% Rank-1 improvement. Finally, the CSSE structural experts boost performance to an optimal of 60.8% mAP and 79.0% Rank-1. These results validate the effectiveness and generalizability of each proposed component.
Table 8: Effectiveness of sampling strategies.
Methods mAP R-1
A All-Token 79.0 83.0
B Top-$K$ 81.4 84.3
C Fixed-$\sigma$ 80.3 85.0
D $\bm{\sigma}_{m}$(Ours) 82.4 86.6
Table 9: Effectiveness of caption quality.
Methods Caption mAP R-1
IDEA IDEA-Text 80.2 82.1
NEXT-Text 80.2 84.0
NEXT IDEA-Text 80.5 84.7
NEXT-Text 82.4 86.6
B.2 Effectiveness of Sampling Strategies
As shown in Table 8, we analyze different semantic expert sampling strategies on the RGBNT201 [14] dataset to evaluate their impact on model performance. In Row A, the All-Token strategy encodes all patch tokens $\mathbf{F}_{I,m}^{\mathrm{tok}}$ for multi-modal fusion, but fails to identify key semantic regions across modalities, resulting in significant performance degradation. Row B adopts the Top-$K$ strategy, replacing the dynamic routing threshold $\bm{\sigma}_{m}$ in $\mathbf{R}_{m}$ with the top 50% tokens, achieving a suboptimal result of 81.4% mAP and 84.3% Rank-1. Row C replaces the dynamically generated threshold from $\mathbf{f}_{I,m}^{\mathrm{cls}}$ with a fixed value (Fixed-$\sigma$), which lacks adaptive perception and relies on handcrafted heuristics, leading to inferior performance. In contrast, our proposed dynamic sampling strategy enables optimal performance, demonstrating its effectiveness in semantic sampling routing.
B.3 Effectiveness of Caption Quality
Table 9 presents an analysis on the RGBNT201 [14] dataset, evaluating the impact of caption quality on model performance. Compared with the captions used in IDEA, our high-quality captions lead to improved Rank-1 accuracy for the IDEA framework. Conversely, applying IDEA’s captions in our NEXT model results in a -1.9% drop in both mAP and Rank-1. These results demonstrate that caption quality plays a critical role in caption-based methods and highlight the effectiveness of our caption generation approach in NEXT.
Appendix C Reliable Caption Generation
Thanks to the powerful semantic generalization capability of Multi-modal Large Language models (MLLMs), we can leverage their out-of-the-box ability to generate comprehensive textual semantic descriptions about the object appearance. However, simple template-based text generation methods [23] face many practical challenges when deal with multi-modal data, such as noise interference from low-quality images and significant style discrepancies in some modalities, making it difficult to obtain clear object textual descriptions. To this end, our proposed approach firstly employs multiple MLLMs to generate a collection of attributes with confidence scores, and secondly we leverage the confidence scores with a multi-modal attributes aggregation strategy to generate clear and comprehensive object identity textual descriptions.
Confidence-aware Attributes Generation. To flexibly analyze the appearance information of the object, we use the generation of identity attributes as the basic step of reliable caption generation. We first define an instruction template to guide multiple MLLMs (e.g., GPT-4o [34], Qwen-VL [35], etc.) in structuring the attribute set of the input image. However, simple structured attribute information cannot quantitatively measure the confidence level of each item, the model suffers from low-quality visual modality noise, which leads to incorrect attribute selection. To address this, we introduce the concept of confidence score in the instruction template (See Template 6 and 7 for the full attribute template.) This allows us to observe the confidence level of each MLLM assigns to its output for each attribute. Additionally, to avoid the model ignoring environment information, we not only structure the output of ‘age’, ‘gender’, ‘upper clothing’, ‘lower clothing’, ‘hairstyle’, ‘footwear’ and other appearance attributes (in the case of pedestrian), but also require the MLLMs to output ‘view’, ‘illumination’, ‘capture time’, and ‘target clarity’ attributes to aware environmental information.
Figure 6: Instruction template for generating pedestrian attributes with confidence.
Figure 7: Instruction template for generating vehicle attributes with confidence.
Multi-Modal Attributes Aggregation. Based on the structured attribute set with confidence score, we are able to further generate reliable captions for the object. However, as shown in Fig. 2, the RGB modal is affected by lighting degradation, causing the MLLMs to fail recognize the ‘backpack’ attribute. In contrast, the thermal infrared modality for the same identity easily reveals additional ‘backpack’ attribute. This leads to the model generating vague and incomplete textual captions such as ‘unknown,’ ‘unclear,’ or ‘not carrying’ based on mono visual modality. To address this problem, we propose a multi-modal attribute aggregation strategy based on confidence scores to preprocess semantic complementary information in different visual modalities.
(1) Multi-Modal Merge and Complement. Specifically, we first use multiple MLLMs to identify object attributes, and for each identified attribute we select the MLLMs with the highest confidence as their final value to improve the accuracy of each attribute within the modality. Furthermore, taking the RGB modality as an example, its modality-specific low-frequency attributes, such as ‘illumination,’ ‘upper clothing color,’ and ‘lower clothing color,’ are easy to recognize. In contrast, attributes requiring high-frequency information, such as ‘carrying,’ ‘holding,’ ‘handbag,’ and ‘backpack,’ are often incorrectly recognized as ‘unknown,’ ‘unclear,’ or ‘not carrying.’ This significantly affects the clarity of the final generated captions. To address this, we rely on the confidence scores to select the highest-confidence attributes from the NIR and TIR modalities, which are unaffected by lighting conditions, in order to recheck and complement the absent information. Similarly, we apply the same complementary strategy to each modality in order to obtain a complete and reliable set of attributes for each modality.
(2) Caption Generation. Building on the above complete attribute set, we define the following instruction template to guide the LLM [51] in combining the structured attribute set from each modality to generate the final text caption (See Template 8 and 9 for the full caption template.) Finally, we have constructed a complete and reliable caption generation pipeline, and we leverage the same process on vehicle datasets to obtain a comprehensive semantic description of the vehicle.
Figure 8: Instruction template for generating pedestrian caption.
Figure 9: Instruction template for generating vehicle caption.
Appendix D Visualization Analysis
D.1 Feature Distribution
As shown in Fig. 10, to further investigate the impact of the proposed modules on model features, we visualize the feature distributions on the RGBNT201 [14] dataset using T-SNE. Comparing Fig. 10 (a) and 10 (b), the introduction of MMFA effectively aggregates different instances of the same identity. As illustrated in Fig. 10 (c), incorporating the TMSE semantic experts improves the separability of samples across identities. In Fig. 10 (d), the inclusion of CSSE structural experts leads to tighter clustering of difficult identity samples, resulting in well-separated intra-class and inter-class distributions. These results validate the effectiveness of each proposed module.
Figure 10: T-SNE [53] visualization of the feature distribution on RGBNT201 [14] (a) Baseline, (b) Baseline + MMFA, (c) Baseline + MMFA + TMSE, and (d) Baseline + MMFA + TMSE + CSSE (Ours).
Figure 11: Visualization of the sampled patch tokens and activated feature regions for experts on RGBNT201 [14].
Figure 12: Visualization of the sampled patch tokens and activated feature regions for experts on MSVR310 [13].
D.2 Expert Feature Visualization
As shown in Fig. 11 and Fig. 12, to further explore the recognition and perception capabilities of each expert, we visualize the sampled tokens of semantic experts and the activated regions of structural experts. Clearly, we observe that each semantic expert $\textbf{E}_{T}=\{E_{t}^{(1)},E_{t}^{(2)},\dots,E_{t}^{(N_{T})}\}$ tends to focus on tokens related to the object, which allows the experts to attend to distinct semantic parts. Moreover, each semantic expert preferentially samples tokens from different spectral modalities. The semantic experts effectively sample complementary tokens across modalities, demonstrating their ability to perceive fine-grained object parts and extract multi-modal features under text modulation. Furthermore, as illustrated in Fig. 11 and Fig. 12, we observe that structural experts $\textbf{E}_{C}=\{E_{c}^{(1)},E_{c}^{(2)},\dots,E_{c}^{(N_{C})}\}$ consistently perceive the complete structure within modalities, maintaining coarse-grained structural features of the object. These visualizations strongly support the effectiveness of NEXT’s multi-grained expert design that models ReID through semantic sampling and structural perception.
D.3 Retrieval Results
To verify the model’s effectiveness in real-world scenarios, we visualize the retrieval results of models with different module combinations on the RGBNT201 [14] and MSVR310 [13] datasets and compare NEXT with state-of-the-art methods [18, 23, 24]. As shown in Fig. 13, on the RGBNT201 [14] dataset, the introduction of different modules enables the model to more effectively recognize pedestrians with the same identity. Fig. 14 compares NEXT with existing state-of-the-art methods, demonstrating that NEXT achieves superior retrieval performance for target pedestrians under various low-light conditions. In Fig. 15, on the MSVR310 [13] dataset, the model’s ability to identify challenging vehicles progressively improves with the addition of different components. Fig. 16 compares NEXT to current state-of-the-art methods, showing that NEXT effectively leverages multi-modal data to overcome challenges in low-light and cross-day-night scenarios, accurately identifying vehicles.
Figure 13: Visualization of the retrieval results on RGBNT201 [14] (a) Baseline, (b) Baseline + MMFA, (c) Baseline + MMFA + TMSE, and (d) Baseline + MMFA + TMSE + CSSE (Ours).
Figure 14: Visualization of the retrieval results on RGBNT201 [14] (a) DeMo [18], (b) IDEA [23], and (c) NEXT (Ours).
Figure 15: Visualization of the retrieval results on MSVR310 [13] (a) Baseline, (b) Baseline + MMFA, (c) Baseline + MMFA + TMSE, and (d) Baseline + MMFA + TMSE + CSSE (Ours).
Figure 16: Visualization of the retrieval results on RGBNT201 [14] (a) DeMo [18], (b) ICPL-ReID [24], and (c) NEXT (Ours).
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The claims in the abstract and introduction are well aligned with the paper’s contributions and overall scope.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitation of our work in Appendix E.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [No]
  14. Justification: This paper is primarily empirical, focusing on model architecture and evaluation.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide detailed experimental settings in Section 4.1 and Appendix A.2.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Our code will be released upon acceptance, and implementation details are provided in Appendix A.2 to assist readers.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The paper includes all necessary information, such as data splits optimizer configuration, and hyperparameter selection, to support reproducibility and result verification in Appendix A.2.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Statistical significance analysis and error bars are not included, as the experiments are deterministic and yield consistent results under fixed conditions.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We include information about the computing environment (CPU, GPU, and memory) in Appendix A.2 to ensure our experiments can be reliably reproduced.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have carefully reviewed the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We provide a discussion of the broad societal impacts in Appendix E.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [No]
  54. Justification: This paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have properly cited and acknowledged relevant prior works.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: This work does not involve the release of any new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This work does not involve crowdsourcing nor research involving human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This work does not involve crowdsourcing nor research involving human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: This study uses LLMs and MLLMs as a component of the research methodology.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.